### v2

@@ 6 Limitations
The fact that Faetar has no standard orthography does not detract from the utility of ASR: a major application of ASR for minoritized and endangered languages is in indexing audio recordings and in generating textual resources from them (that may, for example, later be used to build educational materials), in addition to linguistic research. Document collections can be built without a standard orthography; nevertheless, to be useful, the language community eventually needs to be able to use and understand the transcription system. Lessons learned through this benchmark are also relevant to speech-only applications (speech-to-speech translation, spoken language generation and dialogue), but with the limited amount of data, the feasibility of these traditionally more resource-intensive tasks remains to be seen.
7 Summary of Contributions
We have presented the Faetar Automatic Speech Recognition Benchmark, an open benchmark corpus with evaluation and baseline code aimed at very under-resourced languages. We demonstrate that out-of-the-box application of state-of-the-art multilingual foundation models yields an improvement over spectral features, and that using unlabelled data yields further improvements. Nevertheless, the resulting PER would generally be unacceptable without manual correction. We release the benchmark with the aim of encouraging careful attention to corpora and languages like Faetar, to deepen our commitment to the promise of universal speech technology.

### v3

@@ VI Limitations
The fact that Faetar has no standard orthography does not detract from the utility of ASR: a major application of ASR for minoritized and endangered languages is in indexing audio recordings and in generating textual resources from them (that may, for example, later be used to build educational materials), in addition to linguistic research. Document collections can be built without a standard orthography. Nevertheless, to be useful, the language community eventually needs to be able to use the transcription system.
Lessons learned through this benchmark are also relevant to speech-only applications (speech-to-speech translation, spoken language generation and dialogue), but with this limited amount of data, the feasibility of these traditionally more resource-intensive tasks remains to be seen.
VII Summary of Contributions
We have presented the Faetar Automatic Speech Recognition Benchmark, an open benchmark corpus with evaluation and baseline code aimed at very under-resourced languages. We demonstrate that out-of-the-box application of state-of-the-art multilingual foundation models yields an improvement over spectral features, and that using unlabelled data yields further improvements. The fact that we obtain improvements on the constrained ASR task by varying models suggests that there may be further improvements to be made to the model architectures as well. Nevertheless, in all cases, the baseline PERs would generally be unacceptable without manual correction. We leave additional approaches (e.g., improved speaker modelling, speech enhancement, hybrid models) for the community to explore, and release the benchmark with the aim of encouraging careful attention to corpora and languages like Faetar, deepening the field’s commitment to the promise of universal speech technology.